AIR National’s capacity is 120 passengers per flight. It currently carries 74 passengers per flight. Growth in passengers is expected to be 6 percent annually. New planes will have to be ordered when the company is carrying 90 percent of capacity. How long will it be before the firm must order new planes? PLEASE PROVIDE INPUTS AND WHAT TO SOLVE FOR IN THE FINANCIAL CALCULATOR DO NOT NEED FORMULAS. USING THE BA II PLUS
Answer #1

90 percent of capacity means=120*90%=108 passengers

use financial calculator as below

PV=-74 is the current or present value

FV=108 as the future value

I/Y=6 is the annual rate of growth

PMT=0

Click CPT

Click N=6.49 years

the above is answer..